# Quantum metrology through spectral measurements in quantum optics

### v1

@@ III.1 Spectral dependence of the Fisher information
In Fig. 2(a), we show a frequency slice of $F_{\Delta}$ for $\Omega=3\gamma$ (black solid curve), revealing a structure reminiscent of the Mollow triplet. This behavior reflects the underlying frequency-dependent features of the resonance fluorescence spectrum, characterized by the steady-state sensor population [42, 43, 44], $\langle\hat{\xi}^{\dagger}\hat{\xi}\rangle\sim\text{Re}\int_{0}^{\infty}d\tau e^{i\Delta_{\xi}\tau}\langle\hat{\sigma}^{\dagger}(0)\hat{\sigma}(\tau)\rangle$. Prominent features in $F_{\Delta}$ appear near the resonant frequencies of the Mollow triplet: $F_{\Delta}\rightarrow 0$ ($\Delta_{\xi}\approx 0$) and two local maxima ($\pm 2\Omega$). These features indicate a well-defined spectral structure in terms of the sensor-laser detuning $\Delta_{\xi}$, revealing frequency windows where the estimation protocol is either significantly enhanced or strongly suppressed. However, additional nontrivial features emerge that are not captured by the first-order coherence alone, such as, the local minima and maxima near $\Delta_{\xi}\approx\pm\Omega$. In fact, these features are more closely related to higher-order moments, $\langle\hat{\xi}^{\dagger k}\hat{\xi}^{k}\rangle$ with $k\geq 2$, which reflect multi-photon processes beyond the fluorescence spectrum. This connection becomes evident when comparing $F_{\Delta}$ to higher-order correlators, such as $\langle\hat{\xi}^{\dagger 2}\hat{\xi}^{2}\rangle$, $\langle\hat{\xi}^{\dagger 3}\hat{\xi}^{3}\rangle$, and $\langle\hat{\xi}^{\dagger 4}\hat{\xi}^{4}\rangle$, which show qualitative similarities with the additional structure in the Fisher information, as shown in the red dashed curves in Fig. 2(a).
These observations highlight a central point: although the CFI in Eq. (5) depends solely on the diagonal elements of the sensor density matrix, it cannot be straightforwardly interpreted in terms of standard optical observables. The underlying reason is that the photon-counting probability $p(n|\theta)$ [see Eq. (6)] is inherently a sum of nonlinear functions of the normally ordered moments of the field. Specifically, they can be expressed as an infinite sum over normally ordered correlators [60, 86],
@@ III.2 Impact of sensor linewidth in parameter estimation
In our setup, each sensor acts as a frequency filter, modeled by a Lorentzian band-pass filter function with linewidth $\Gamma$ [42, 43, 44], providing a faithful physical description of photo-detection [41]. The linewidth $\Gamma$ sets the spectral ($\Delta\omega\sim\Gamma$) and temporal ($\Delta t\sim 1/\Gamma$) resolution of the measurement, as required by the Heisenberg principle [41, 87]. The impact of spectral filtering has been extensively studied in the context of a coherently driven qubit, where depending on the sensor linewidth the photon correlations [49, 50, 51, 52], and the degree of photonic entanglement  [53, 40], can be tuned as the effective description of the photonic state in the sensor is influenced by the spectral resolution [88]. Therefore, given its central role in the detection process and the properties defining the quantum photonic state, we now assess the metrological impact of the sensor linewidth.
@@
Figure 3 shows the CFI as a function of the sensor linewidth $\Gamma$ and driving strength $\Omega$ for a fixed value of the sensor frequency, $\Delta_{\xi}=\Omega$. Figure 3(a) depicts the CFI in terms of $\Gamma$ for two particular choices of the driving strength: the Heitler regime ($\Omega=\gamma/2\sqrt{2}$, blue solid) and the Mollow triplet ($\Omega=3\gamma$, blue dashed). Both regimes feature an optimal detection window that maximizes the CFI, roughly in the range $\Gamma_{\text{opt}}\in(10^{-2}\,\Omega,10\,\Omega)$. Generally, in both extreme limits of the sensor linewidth—narrowband ($\Gamma\rightarrow 0$) and broadband ($\Gamma\rightarrow\infty$)—the CFI reaches minimal values. In the narrowband regime, the sensor achieves perfect spectral resolution but loses time definition, leading to a Gaussian description of the signal. As a result, when $\Gamma\ll 10^{-2}\Omega$, the CFI collapses onto the SNR (red curves), $F_{\theta}\approx\text{SNR}_{\theta}$, indicating that estimation is dominated by mean values. This behavior is consistent with previous reports observing frequency-resolved correlations at optimal ranges of filter bandwidth, narrow enough to resolve spectral features, but broad enough to limit the time uncertainty in the detection [49, 50, 53, 40].
@@
The comparison between $F_{\theta}$ and the SNR provides some intuition about the importance of quantum fluctuations in parameter estimation. An alternative measure for this assessment, the Poissonian Fisher information (PFI), was introduced by the authors in Ref. [89]. The PFI is based on a Poissonian approximation for the photon statistics [90], leading to an expression of the Fisher information given by [see Appendix E],
@@
Unlike the SNR, the PFI assumes coherent-state statistics, where the variance equals the mean. In regimes where the signal is Gaussian but not fully coherent the this quantity may misestimate the accessible information. Nevertheless, $F^{P}_{\theta}\approx\text{SNR}_{\theta}$ when $\langle\Delta^{2}\hat{\xi}^{\dagger}\hat{\xi}\rangle\approx\langle\hat{\xi}^{\dagger}\hat{\xi}\rangle$, as shown by the purple dot-dashed curves in Fig. 3(a). Therefore, for faithful assessment of estimation performance when only the mean value is accessible, the SNR provides a more reliable benchmark than $F^{P}_{\theta}$, as it properly accounts for the actual statistical properties of the signal.
In the broadband limit, the measurement becomes color-blind: temporal precision is perfectly defined, but spectral information is lost. As a result, the CFI decreases and becomes equal to the SNR, signaling a reduction of the information contained in quantum fluctuations of the filtered field.
These results showcase some of the nuances involved in the process of filtering radiation for the task of parameter estimation, as well as the potential improvements one can achieve by optimizing the process. When operating in the optimal detection regime [see Fig. 3(b)], the sensor can boost the CFI by over ten orders of magnitude. Nonetheless, even in regimes where the CFI, SNR, and PFI get drastically supressed—due to either extreme spectral or time resolution—, the quantum Fisher information (QFI) remains finite, revealing the existence of an optimal POVM different from the standard photon-counting measurement used here. Below, we discuss an update of the measurement scheme that can bring us closer to the limit set by the QFI.
Figure 3: Impact of sensor linewidth $\Gamma$ on the performance of the estimation protocol for a single sensor. (a) $I_{\Delta}$ (black), $F_{\Delta}$ (blue), and $\text{SNR}_{\Delta}$ (red) as a function of $\Gamma$, shown for two values of $\Omega=\{1/2\sqrt{2},3\}\gamma$, represented by solid and dashed curves, respectively. (b) $F_{\Delta}$ in terms of the driving strength $\Omega$ and the sensor linewidth $\Gamma$. The optimal regime is approximately bounded by $\Gamma\in(10^{-2},10)\Omega$, i.e., where the sensor preserves spectral resolution and metrological performance is enhanced. Parameters: $\Delta=10^{-6}\gamma,\ \Delta_{\xi}=\Omega,\ \varepsilon=1$.
@@ III.3 Mean field engineering
Figure 4: Optimizing quantum parameter estimation via mean-field engineering for a single sensor. (a) $F^{\alpha}_{\Delta}$ over the phase space of the local oscillator displacement. Two points are highlighted: the displacement $\alpha_{\text{fluct}}\approx-1.58i$ that isolates the quantum fluctuations, and the optimal displacement $\alpha_{\text{opt}}\approx 3.95-1.56i$ that maximizes the Fisher information. The inset illustrates the mean-field engineering scheme, where the signal $\hat{s}$ is mixed with a coherent field $\alpha_{0}$ via an unbalanced beam splitter in the limit of perfect transmittance, yielding an engineered signal $\hat{s}+\alpha$. (b) $I_{\Delta}$ (gray solid), $F_{\Delta}$ (blue solid), $F_{\Delta}^{\alpha_{\text{fluct}}}$ (blue dot-dashed), and $F_{\Delta}^{\alpha_{\text{opt}}}$ (blue dashed) in the frequency domain ($\Delta_{\xi}$). (c) Same quantities as in (b), now plotted against the emitter-laser detuning ($\Delta$), with $\Delta_{\xi}=\Omega$. Parameters: (a-c) $\Omega=\gamma/2\sqrt{\gamma}$, $\Gamma=10^{-1}\gamma$, $\varepsilon=0.5$; (a,b) $\Delta=10^{-6}\gamma$; (a) $\Delta_{\xi}=0$; (c) $\Delta_{\xi}=\Omega$.
A recent experimental work [72] demonstrated that adjusting the weight of the coherent component of light via interference [91, 92, 93, 94] serves as a novel control knob for engineering quantum statistics, termed mean-field engineering. This was demonstrated by unlocking multiphoton effects at very low-driving in resonance fluorescence. While the bare emission is antibunched, the engineered emission becomes superbunched at all orders when the classical field is perfectly suppressed. The nature of the multiphoton emission can thus be tuned by adjusting the amplitude of the interfering classical field [88, 71, 72]. The potential of this mean-field engineering technique in quantum metrology and frequency-resolved detection is still unexplored.
@@
Figure 4(a) shows the CFI s a function of the displacement $\alpha$ in phase space at a fixed frequency $\Delta_{\xi}$, evidencing that displacing the signal away from $\alpha=0$—which corresponds to standard photon-counting, illustrated by the blue solid curve in Fig. 4(b)—can significantly enhance the CFI. Following the idea presented in Ref. [71, 72], we first consider the displacement that completely cancels the coherent component and isolates quantum fluctuations by tuning the amplitude to (see Appendix F):
@@
shown as a point in Fig. 4(a) and across the frequency domain in Fig. 4(b), denoted by $F_{\Delta}^{\alpha_{\text{fluct}}}$ (dot-dashed). Complete cancellation of the coherent component modifies the metrological performance, but it does not always results in an improvement; the effect depends on the spectral region: e.g., it improves sensitivity near resonance but not at $\Delta_{\xi}\approx\Omega$.
More generally, we can optimize the measurement by choosing the displacement $\alpha_{\text{opt}}\equiv\arg[\max_{\alpha}F_{\theta}^{\alpha}]$ that yields the maximum CFI, $F_{\theta}^{\alpha_{\text{opt}}}$. The optimum displacement is also indicated as a point in Fig. 4(a). This protocol significantly enhances the CFI and can even saturate the quantum Fisher information across the full frequency range, as shown in Fig. 4(b).
Further information on the metrological impact of mean-field engineering is shown in Fig. 4(c), where we compare the QFI and the CFIs across values of the estimated parameter, $\Delta$, for the three relevant displacements mentioned here: $\alpha=0$ (standard photon-counting), $\alpha_{\text{fluct}}$ (isolation of fluctuations), and $\alpha_{\text{opt}}$ (optimal measurement). All curves show a global maximum near $\Delta\approx\Omega$, explicitly showing that the most sensitive point is not the resonant case, as it was already reported in Ref. [34]. As in panel (b), the optimally displaced CFI provides the maximum achievable information that saturates the QFI, with only minor deviations near resonance ($\Delta=0$), thereby reaching the ultimate metrological limit.
These results establish mean-field engineering in combination of frequency-resolved photo-detection as a powerful method to maximize parameter sensitivity.

### v2

@@ III.1 Spectral dependence of the Fisher information
In Fig. 2(a), we show a frequency slice of $F_{\Delta}$ for $\Omega=3\gamma$ (black solid curve), revealing a structure reminiscent of the Mollow triplet. This behavior reflects the underlying frequency-dependent features of the resonance fluorescence spectrum, characterized by the steady-state sensor population [44, 45, 46], $\langle\hat{\xi}^{\dagger}\hat{\xi}\rangle\sim\text{Re}\int_{0}^{\infty}d\tau e^{(i\Delta_{\xi}-\Gamma/2)\tau}\langle\hat{\sigma}^{\dagger}(0)\hat{\sigma}(\tau)\rangle$. Prominent features in $F_{\Delta}$ appear near the resonant frequencies of the Mollow triplet: $F_{\Delta}\rightarrow 0$ ($\Delta_{\xi}\approx 0$) and two local maxima ($\pm 2\Omega$). These features indicate a well-defined spectral structure in terms of the sensor-laser detuning $\Delta_{\xi}$, revealing frequency windows where the estimation protocol is either significantly enhanced or strongly suppressed. However, additional nontrivial features emerge that are not captured by the first-order coherence alone, such as, the local minima and maxima near $\Delta_{\xi}\approx\pm\Omega$. In fact, these features are more closely related to higher-order moments, $\langle\hat{\xi}^{\dagger k}\hat{\xi}^{k}\rangle$ with $k\geq 2$, which reflect multi-photon processes beyond the fluorescence spectrum. This connection becomes evident when comparing $F_{\Delta}$ to higher-order correlators, such as $\langle\hat{\xi}^{\dagger 2}\hat{\xi}^{2}\rangle$, $\langle\hat{\xi}^{\dagger 3}\hat{\xi}^{3}\rangle$, and $\langle\hat{\xi}^{\dagger 4}\hat{\xi}^{4}\rangle$, which show qualitative similarities with the additional structure in the Fisher information, as shown in the red dashed curves in Fig. 2(a).
These observations highlight a central point: although the CFI in Eq. (8) depends solely on the diagonal elements of the sensor density matrix, it cannot be straightforwardly interpreted in terms of standard optical observables. The underlying reason is that the photon-counting probability $p(n|\theta)$ [see Eq. (9)] can be expressed as an infinite sum over normally ordered correlators [48, 95],
@@ III.2 Impact of sensor linewidth in parameter estimation
In our setup, each ancillary sensor acts as a frequency filter—see Eq. (4)—, modeled by a Lorentzian band-pass filter function with linewidth $\Gamma$ [44, 45, 46], providing a faithful physical description of photo-detection [42]. The linewidth $\Gamma$ sets the spectral ($\Delta\omega\sim\Gamma$) and temporal ($\Delta t\sim 1/\Gamma$) resolution of the measurement, as required by the uncertainty relations between energy and time [42, 96]. The impact of filter linewidth on photonic statistics has been extensively studied in the context of a coherently driven qubit. In particular, the time–energy uncertainty introduced by filtering has a profound effect on frequency-resolved correlations [54, 55, 56, 97, 57]. While color-blind measurements yield perfect antibunching, filtering introduces a finite time uncertainty that lifts this perfect single-photon character and reveals a rich landscape of frequency-resolved correlations. This non-trivial dependence extends to other quantities as well, including the degree of photonic entanglement [58, 41]. Given its central role in the detection process and the properties defining the quantum photonic state, we now assess the metrological impact of the sensor linewidth.
@@
Figure 3 shows the CFI as a function of the sensor linewidth $\Gamma$ and driving strength $\Omega$ for a fixed value of the sensor frequency, $\Delta_{\xi}=\Omega$. Figure 3(a) depicts the CFI in terms of $\Gamma$ for a particular choice of the driving strength ($\Omega=3\gamma$, blue solid). Although this value lies in the Mollow regime, we find that both the Heitler and Mollow regimes exhibit an optimal detection window that maximizes the CFI, occurring roughly in the range $\Gamma_{\text{opt}}\approx(10^{-2}\,\Omega,10\,\Omega)$. Generally, in both extreme limits of the sensor linewidth—narrowband ($\Gamma\rightarrow 0$) and broadband ($\Gamma\rightarrow\infty$)—the CFI reaches minimal values. In the narrowband regime, the sensor achieves perfect spectral resolution but loses time definition, leading to a Gaussian description of the signal. As a result, when $\Gamma\ll 10^{-2}\Omega$, the CFI collapses onto the SNR (red curve), $F_{\theta}\approx\text{SNR}_{\theta}$, indicating that estimation is dominated by mean values. This behavior is consistent with previous reports observing frequency-resolved correlations at optimal ranges of filter bandwidth, narrow enough to resolve spectral features, but broad enough to limit the time uncertainty in the detection [54, 55, 58, 41].
Figure 3: Impact of sensor linewidth $\Gamma$ on the performance of the estimation protocol for a single sensor. (a) $I_{\Delta}$ (grey solid), $F_{\Delta}$ (blue solid), and $\text{SNR}_{\Delta}$ (red solid) as a function of $\Gamma$, shown for $\Omega=3\gamma$. Additionally, $F_{\Delta}^{\alpha_{\text{opt}}}$ (blue dashed) and $F_{\Delta}^{P}$ (yellow dot-dashed) are shown. (b) $F_{\Delta}$ in terms of the driving strength $\Omega$ and the sensor linewidth $\Gamma$. The optimal regime is approximately bounded by $\Gamma\approx(10^{-2}\Omega,10\Omega)$, i.e., where the sensor preserves spectral resolution and metrological performance is enhanced. Parameters: $\Delta=10^{-6}\gamma,\ \Delta_{\xi}=\Omega,\ \varepsilon=1$.
@@
The comparison between $F_{\theta}$ and $\text{SNR}_{\theta}$ provides some intuition about the importance of quantum fluctuations in parameter estimation. An alternative measure for this assessment, the Poissonian Fisher information (PFI), was introduced by the authors in Ref. [98]. The PFI is based on a Poissonian approximation for the photon statistics [99], leading to an expression of the Fisher information given by [see Appendix E],
@@
Unlike the SNR, the PFI assumes coherent-state statistics, where the variance equals the mean. In regimes where the signal is Gaussian but not fully coherent the this quantity may misestimate the accessible information. Nevertheless, $F^{P}_{\theta}\approx\text{SNR}_{\theta}$ when $\langle\Delta^{2}\hat{\xi}^{\dagger}\hat{\xi}\rangle\approx\langle\hat{\xi}^{\dagger}\hat{\xi}\rangle$, as shown by the yellow dot-dashed curve in Fig. 3(a). Therefore, for faithful assessment of estimation performance when only the mean value is accessible, the SNR provides a more reliable benchmark than $F^{P}_{\theta}$, as it properly accounts for the actual statistical properties of the signal.
In the broadband limit, the measurement becomes color-blind: temporal precision is perfectly defined, but spectral information is lost. As a result, the CFI decreases and becomes equal to the SNR—and to the PFI—, signaling a reduction of the information contained in quantum fluctuations of the filtered field.
These results showcase some of the nuances involved in the process of filtering radiation for the task of parameter estimation, as well as the potential improvements one can achieve by optimizing the process. When operating in the optimal detection regime [see Fig. 3(b)], the sensor can boost the CFI by over ten orders of magnitude. Nonetheless, even in regimes where the CFI, SNR, and PFI get drastically suppressed—due to either extreme spectral or time resolution—, the quantum Fisher information (QFI) remains finite, revealing the existence of an optimal POVM different from the standard photon-counting measurement used here. Below, we discuss an update of the measurement scheme that can bring us closer to the limit set by the QFI.
@@ III.3 Mean field engineering
Recent experimental works [80, 81] showed that adjusting the weight of the coherent component of light via interference [100, 101, 102, 103] serves as a novel control knob for engineering quantum statistics, termed mean-field engineering. This was verified by unlocking multiphoton effects at very low-driving in resonance fluorescence. While the bare emission is antibunched, the engineered emission becomes superbunched at all orders when the classical field is perfectly suppressed. The nature of the multiphoton emission can thus be tuned by adjusting the amplitude of the interfering classical field [97, 79, 80, 81]. The potential of this mean-field engineering technique in quantum metrology and frequency-resolved detection is still unexplored.
@@
Figure 4: Optimizing quantum parameter estimation via mean-field engineering for a single sensor. (a) Mean-field engineering scheme: the multi-mode radiation is mixed with a coherent field via an unbalanced beam splitter in the limit of perfect transmittance, yielding an engineered signal that is frequency-filtered. (b) $F^{\alpha}_{\Delta}$ over the phase space of the local oscillator displacement. Two points are highlighted: the displacement $\alpha_{\text{fluct}}\approx-1.58i$, and the optimal displacement $\alpha_{\text{opt}}\approx-4.3-1.4i$. (c) $I_{\Delta}$ (grey solid), $F_{\Delta}$ (blue solid), $F_{\Delta}^{\alpha_{\text{fluct}}}$ (blue dot-dashed), and $F_{\Delta}^{\alpha_{\text{opt}}}$ (blue dashed) in the frequency domain ($\Delta_{\xi}$). (d) Same quantities as in (c), now plotted against the emitter-laser detuning ($\Delta$), with $\Delta_{\xi}=\Omega$. Parameters: (b-d) $\Omega=\gamma/2\sqrt{\gamma}$, $\Gamma=10^{-1}\gamma$, $\varepsilon=0.5$; (b,c) $\Delta=10^{-6}\gamma$; (b) $\Delta_{\xi}=0$; (d) $\Delta_{\xi}=\Omega$.
Figure 4(b) shows the CFI as a function of the displacement $\alpha$ in phase space at a fixed frequency $\Delta_{\xi}$, evidencing that displacing the signal away from $\alpha=0$—which corresponds to standard photon-counting, illustrated by the blue solid curve in Fig. 4(c)—can significantly enhance the CFI. Following the idea presented in Ref. [79, 80, 81], we first consider the displacement that completely cancels the coherent component and isolates quantum fluctuations by tuning the amplitude to (see Appendix F):
@@
shown as a point in Fig. 4(b) and across the frequency domain in Fig. 4(c), denoted by $F_{\Delta}^{\alpha_{\text{fluct}}}$ (dot-dashed). Complete cancellation of the coherent component modifies the metrological performance, but it does not always result in an improvement; the effect depends on the spectral region: e.g., it improves sensitivity near resonance but not at $\Delta_{\xi}\approx\Omega$.
More generally, we can optimize the measurement by choosing the displacement $\alpha_{\text{opt}}\equiv\arg[\max_{\alpha}F_{\theta}^{\alpha}]$ that yields the maximum CFI, $F_{\theta}^{\alpha_{\text{opt}}}$. The optimum displacement is also indicated as a point in Fig. 4(b). This protocol significantly enhances the CFI and can even saturate the quantum Fisher information across the full frequency range, $F_{\theta}^{\alpha_{\text{opt}}}\approx I_{\theta}$, as shown in Fig. 4(c). Intuitively, interfering the signal with a tunable local oscillator effectively reshapes how the parameter dependence is distributed among the normally ordered moments that enter the CFI [see Eq. (16)]. By adjusting the displacement, one can enhance the relative change of these moments with respect to the parameter $\theta$, thereby altering—and potentially increasing—the achievable Fisher information. We emphasize that the local oscillator does not introduce new information; rather, the displacement modifies how the existing parameter dependence of $\hat{\rho}_{\theta}$ and $\partial_{\theta}\hat{\rho}_{\theta}$ appears in the normally ordered moments that determine the CFI. For this reason, neither $\alpha=0$ nor $\alpha_{\mathrm{fluct}}$ correspond to privileged operating points: they are simply two particular choices within a continuous set of possible displacements. The optimal displacement $\alpha_{\mathrm{opt}}$ emerges from fully exploiting this phase-space freedom to maximize the Fisher information, as illustrated in Fig. 4(b).
Further information on the metrological impact of mean-field engineering is shown in Fig. 4(d), where we compare the QFI and the CFIs across values of the estimated parameter, $\theta=\Delta$, for the three relevant displacements mentioned here: $\alpha=0$ (standard photon-counting), $\alpha_{\text{fluct}}$ (isolation of fluctuations), and $\alpha_{\text{opt}}$ (optimal measurement). All curves show a global maximum near $\Delta\approx\Omega$, explicitly showing that the most sensitive point is not the resonant case, as it was already reported in Ref. [35]. As in panel (c), the optimally displaced CFI provides the maximum achievable information that saturates the QFI, with only minor deviations near resonance ($\Delta=0$), thereby reaching the ultimate metrological limit.
These results establish mean-field engineering, combined with frequency-resolved photo-detection, as a powerful method to maximize parameter sensitivity. In fact, beyond the examples presented in Fig. 4, we confirm the validity of this technique by optimizing the performance shown earlier in Fig. 3(a), where the optimized CFI (blue dashed curve) saturates the QFI across the full range of $\Gamma$. As we will see later, this strategy naturally extends to the multi-mode case, where one optimizes a coherent displacement for each filtered mode.
We also observe that the QFI does not inherit the intricate spectral structure visible in the CFI, but instead exhibits a smooth, broad shape centered at the drive resonance. Intuitively, the optimal measurement extracts information most efficiently near resonance, where the collected mode overlaps maximally with the emitted radiation carrying information about $\theta$. Moving away from resonance reduces this overlap and decreases the achievable precision. From a theoretical perspective, displaced photon counting does not span the full POVM space of the radiation, though varying the coherent displacement already explores a large measurement manifold that might contain the optimal measurement. Therefore, optimizing the CFI over $\alpha$ effectively recombines contributions from all correlator orders in Eq. (16), suppressing cancellations and yielding smoother variations in parameter space.